# Fully Parallelized BP Decoding for Quantum LDPC Codes Can Outperform BP-OSD

### v2

@@ 5.3. Complexity and Parameter Selection
Next, we analyze the computational complexity of the proposed decoder. To ensure a fair comparison with baseline methods, we measure the average number of BP iterations under a serial execution model. Specifically, when the initial BP decoding fails, each test vector is decoded sequentially, and the total number of iterations is defined as the cumulative number of BP iterations required until the first successful decoding. This approach provides a conservative estimate of decoding cost, as it does not account for the inherent parallelism of our method, but allows for a meaningful comparison with standard decoders.
@@
Figure 9 shows the growth in decoding complexity for the $\llbracket 144,12,12\rrbracket$ code as we target progressively lower logical error rates. For the BP decoder, we vary the maximum number of iterations to control decoding complexity. For the proposed decoder, we fix the maximum number of iterations per BP instance to 100 and vary $n_{s}$ (the number of test vectors sampled per weight), while keeping $w_{\text{max}}$ constant. This allows us to explore the trade-off between complexity and performance. The physical error rate is fixed at $3\times 10^{-3}$ under the circuit-level noise model. All data points are collected by simulating 10,000 shots for logical error rates above $10^{-3}$ and 100,000 shots for those below.
Across all decoders, we observe a linear region in which the number of BP iterations increases approximately linearly as the logical error rate decreases, up to a point where each curve drops off sharply, forming what we refer to as a cliff. This cliff marks a regime where the decoder can no longer reliably suppress logical errors within reasonable iteration limits. The proposed decoder consistently postpones this cliff compared to baseline BP, maintaining a lower iteration count at comparable logical error rates. Moreover, increasing $w_{max}$ increases the complexity but extends the linear region further and delays the start of the cliff, providing a tunable trade-off between decoding complexity and error suppression.
@@
Figure 9. Complexity growth of different decoders. The parameter $|\Phi|$ is set to 50 for all proposed decoders. The number of iterations is calculated assuming serial execution.
6. Conclusion and Future Work
@@
We introduced a fully parallelizable decoder based on belief propagation (BP). By leveraging speculative decoding and bit-flipping strategies guided by BP oscillation statistics, the proposed method achieves logical error rates comparable to BP-OSD, while significantly reducing computational complexity and avoiding costly Gaussian elimination. Extensive simulations show that our decoder performs exceptionally well under the code-capacity noise model across a range of bivariate bicycle codes. Under the more realistic circuit-level noise model, the decoder still delivers reasonable performance, though it requires a larger number of decoding trials to achieve comparable accuracy.
@@
In future work, we aim to better understand the challenges posed by circuit-level noise and explore targeted improvements, such as more effective candidate selection, improved test vector sampling strategies, efficient decoder implementation, and enhancements to the inner BP decoder, in order to further improve decoding performance in practical fault-tolerant quantum computing systems.
@@ Acknowledgment
The authors would like to thank John Stack for providing source code and for discussions on circuit-level noise simulation, and Timo Hillmann for valuable discussions regarding the BP-OSD decoder. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830. This work was also supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, MPS-2120757, CISE-2217020, and CISE-2316201 as well as DOE DE-SC0025384.
Appendix A Code Construction
In this section, we provide the construction parameters for the qLDPC codes referenced throughout the paper.
A.1. GB codes
@@
(12) $S_{l}=I_{l}>>1,$
where “>>much-greater-than$>$ >” denotes the right cyclic shift for each row in the matrix. For example,
(13) $S_{3}=I_{3}>>1=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}.$
Let $x=S_{l}$ here, The GB code can be defined by two polynomials, $a(x)$ and $b(x)$. These two polynomials can be represented by two matrices $A,B$ naturally, and the parity check matrices for the GB code are defined as
(14) $\begin{split}H_{X}&=[A|B]\\ H_{Z}&=[B^{T}|A^{T}].\end{split}$
The $\llbracket 254,28\rrbracket$ GB code used in this paper can be constructed by: $l=127$, $a(x)=1+x^{15}+x^{20}+x^{28}+x^{66}$, $b(x)=1+x^{58}+x^{59}+x^{100}+x^{121}$ as proposed in (Panteleev and Kalachev, 2021).
A.2. BB Codes
BB codes are constructed similarly to GB codes but with two variables. Let $x=S_{l}\otimes I_{m}$ and $y=I_{l}\otimes S_{m}$, where $\otimes$ denote Kronecker product, the BB codes can also be defined by two polynomials, $A=a(x,y)$ and $B=b(x,y)$. The parity check matrices for BB code are defined similarly as in GB codes.
The BB codes we used in the paper were proposed in (Bravyi et al., 2024). And the polynomials are shown in Table 1.
Table 1. BB Codes Used in Simulations
@@ A.2. BB Codes
A.3. Coprime-BB Codes
The coprime-BB code we used were found in (Wang and Mueller, 2024). Let $\pi=xy$, where $x$ and $y$ are defined the same as in BB codes. The coprime-BB code we tested can be constructed as shown in Table 2.
Table 2. Coprime-BB Codes Used in Simulations
@@ A.3. Coprime-BB Codes
This section presents several codes that demonstrate good performance under BP decoding, along with their corresponding logical error rates. Since both BP-OSD and the proposed decoder act as post-processing techniques, they are only invoked when the BP decoder fails to converge. As a result, for the codes shown below, where BP alone achieves high success rates, we observe similar overall performance across all decoding strategies.
B.1. Code Capacity Model
@@
Figure. 10(a-d) shows the performance of various decoders on several codes under the code-capacity noise model. In these cases, the baseline BP decoder already achieves logical error rates comparable to those of the BP-OSD decoder. In these cases, post-processing yields only marginal improvements.
Figure 10. The logical error rates of different codes under code capacity error model. For (a): $\llbracket 72,12,6\rrbracket$ BB code, BP and BP-OSD have the same error when we set the seed to the same.
B.2. Circuit Level Noise Model
For the $\llbracket 72,12,6\rrbracket$, the BP and BP-OSD decoder have similar performance, as shown in Fig. 11.
Figure 11. Error rate comparison on the $\llbracket 72,12,6\rrbracket$ under circuit-level noise model.

### v3

@@ V-C Complexity and Parameter Selection
Next, we analyze the computational complexity of our BP-SF decoder. To ensure a fair comparison with baseline methods, we measure the average number of BP iterations under a serial execution model. Specifically, when the initial BP decoding fails, each trial syndrome is decoded sequentially, and the total number of iterations is defined as the cumulative number of BP iterations required until the first successful decoding. This approach provides a conservative estimate of decoding cost, as it does not account for the inherent parallelism of our method, but allows for a meaningful comparison with standard decoders.
@@
Fig. 12 shows the growth in decoding complexity for the $\llbracket 144,12,12\rrbracket$ code as we target progressively lower logical error rates. For the BP decoder, we vary the maximum number of iterations to control decoding complexity. For our BP-SF decoder, we fix the maximum number of iterations per BP instance to 100 and vary $n_{s}$ (the number of trial vectors sampled per weight), while keeping $w_{\text{max}}$ constant. This allows us to explore the trade-off between complexity and performance. The physical error rate is fixed at $3\times 10^{-3}$ under the circuit-level noise model. All data points are collected by simulating 10,000 shots for logical error rates above $10^{-3}$ and 100,000 shots for those below.
@@
Figure 12: Complexity growth of different decoders. The parameter $|\Phi|$ is set to 50 for all BP-SF decoders. The number of iterations is calculated assuming serial execution.
Across all decoders, we observe a linear region in which the number of BP iterations increases approximately linearly as the logical error rate decreases, up to a point where each curve drops off sharply, forming what we refer to as a cliff. This cliff marks a regime where the decoder can no longer reliably suppress logical errors within reasonable iteration limits. Our BP-SF decoder consistently postpones this cliff compared to baseline BP, maintaining a lower iteration count at comparable logical error rates. Moreover, increasing $w_{max}$ increases the complexity but extends the linear region further and delays the start of the cliff, providing a tunable trade-off between decoding complexity and error suppression.
Fig. 13 shows how the decoding latency scales as the code size increases. Here, “number of error mechanisms” refers to the total number of error sources in a memory experiment circuit, which grows much faster than the number of qubits. The proposed BP-SF consistently achieves lower average decoding latency. Since most errors are corrected by the initial BP stage, the asymptotic advantage of BP-SF is not evident from the average decoding time alone, which is about $0.63\times$ that of BP-OSD for the $\llbracket 288,12,18\rrbracket$ code. However, when considering only cases where the initial BP fails, BP-SF requires merely $0.1\times$ the latency of BP-OSD, an order of magnitude improvement. Both the overall average latency and the post-processing latency of BP-SF can be further reduced (discussed in Section VI).
Figure 13: Latency scaling of BP-SF and BP-OSD at a physical error rate of $3\times 10^{-3}$. Dashed lines indicate the average latency of the post-processing stage, measured only for cases where the initial BP decoding fails.
VI Parallel Implementation and Performance
In this section, we describe the implementation of our proposed algorithm and compare it with the BP-OSD decoder in the LDPC library [23] and Nvidia’s CUDA-Q library [26]. For a fair comparison, we implemented three versions of our algorithm:
Serial CPU version: We modify the BP decoder from [23], which is also a serial CPU version implemented in C++ and wrapped with Cython, to track bit-level oscillations. If the first decoding attempt fails, trial vectors and corresponding syndromes will be prepared in Python, and each trial syndrome will be decoded using the C++ BP decoder. We return the first successful decoding result and skip the rest.
Parallel CPU version (P=# of worker processes): This version maintains a persistent worker process pool along with input and output queues to handle the remaining $n_{s}\times w_{\max}$ decoding trials after the initial BP attempt fails. The manager process selects candidate positions, generates trial vectors, and computes the corresponding trial syndromes. These trial syndromes are then split into small batches and placed in the input queue. Worker processes continuously retrieve batches from the queue and attempt to decode them until one finds a valid solution, which it places in the output queue. The main process monitors the output queue and, once a valid result is found, signals all workers to stop. The workers then wait for the next input from input queue, and the main process returns the successful result. To avoid accepting stale results, each syndrome is tagged with a serial number. The main process compares the serial number of each fetched result with that of the currently processed syndrome to ensure correctness.
Estimated GPU version (GPU_Est): Since the CUDA-Q library does not provide access to oscillation statistics during decoding, we implemented this version to estimate the decoding speed. We first precompute the syndromes and their corresponding oscillation bits using a CPU-based BP decoder. These syndromes are then decoded using CUDA-Q. If decoding fails, the precomputed oscillation bits are used in Python to generate trial syndromes, which are then decoded one-by-one using CUDA-Q until a successful decoding is found or the maximum number of trials is reached. Due to numerical instability and implementation differences, the CUDA-Q decoder may report a different success status than the CPU BP decoder. In such cases, the pre-computation is invalid and we discard those results. This approach results in a pessimistic estimate of GPU performance, as trial syndrome generation is performed on the CPU after initial failure, and transferring them to the GPU introduces additional memory copy overhead. A more efficient strategy would submit all trial syndromes to the GPU decoder simultaneously and return upon the first success. However, the current CUDA-Q library only supports the decode_batch method, which waits for all syndromes in the batch to complete, effectively blocking on the slowest one.
For the sake of fairness, we select the number of BP iterations to make BP-OSD as fast as possible. At first glance, one might expect that reducing the number of BP iterations would lower the overall decoding latency. However, beyond some point, reducing the number of BP iterations not only degrades the logical error rate but can also increase the total latency as shown in Table I. This is caused by the BP stage being relatively inexpensive compared to the OSD stage, where higher number of BP iterations allows BP to correct more errors reduces the frequency with which the costly OSD procedure must be invoked. We listed the BP iterations of experiments with the $\llbracket 144,12,12\rrbracket$ code under circuit-level noise of $p=3\times 10^{-3}$.
Decoder LER/$d$ @ 3e-3 Avg Time @ 3e-3
BP100-OSD10 2.89 $\times 10^{-4}$ 56.13 ms
BP400-OSD10 2.23 $\times 10^{-4}$ 37.69 ms
BP1000-OSD10 2.11 $\times 10^{-4}$ 36.44 ms
BP2000-OSD10 2.00 $\times 10^{-4}$ 44.01 ms
BP10000-OSD10 1.84 $\times 10^{-4}$ 94.94 ms
TABLE I: Logical error rate per round and average decoding time for BP-OSD with different iterations.
We note that for CPU-based decoders, it is possible to parallelize decoding using OpenMP within each BP iteration or by decoding a batch of input syndromes in parallel. However, we did not pursue these options because the BP/BP-OSD decoder provided by the LDPC library [23] does not currently support OpenMP acceleration. For fairness and consistency, we instead implement parallelization at a coarser granularity, i.e., by decoding trial syndromes in parallel. This approach may be less efficient than parallelizing BP iterations directly, as it only accelerates cases where the initial decoding attempt fails. However, it helps mitigate the long-tail latency caused by the large number of trials required in such cases. We also do not parallelize across input syndromes, as decoding them sequentially is more aligned with real-world use cases, where syndrome extraction is performed sequentially and syndromes arrive in a streaming fashion.
Benchmarks were conducted on an NVIDIA Tesla V100-SXM2-16GB GPU and an Intel Xeon E5-2698 v4 @ 2.20GHz CPU. For each physical error rate, 20,000 syndromes were tested. Our newly proposed decoder with 100 BP iterations, $w_{\max}=10$, and $n_{s}=10$ achieves nearly identical logical error rates to BP1000-OSD10 ($1,000$ BP and OSD-CS of order $10$) below threshold for the $\llbracket 144,12,12\rrbracket$ code, which allows us to adopt these settings for all evaluations in this section.
[…]
Figure 14: Average decoding time per syndrome for the BP-SF decoders compared with BP1000-OSD10 and BP100, under varying physical error rates. BP100 is performed on the CPU serially as a reference lower bound, as it performs no post-processing.
[…]
Figure 15: Distribution of single-syndrome decoding times for different decoder configurations when physical error rate is 0.003. The horizontal violin plots show the probability density of the time taken for each decoding attempt, plotted on a logarithmic scale. Overlaid gray points represent individual decoding events, with vertical jitter added for clarity. The vertical lines mark the min, median, and max decoding time. Decoding events invoking OSD stage are circled in red.
Figure 16 shows the runtime distribution for the GPU decoders. We observe a similar pattern: while the BP-SF method achieves a lower average runtime (5.47 ms vs. 7.37 ms), its maximum runtime (73.74 ms) is higher than that of the BP-OSD decoder (39.76 ms), due to the serial decoding of trial syndromes. This long-tail latency can be effectively mitigated by enabling a GPU function that accepts a batch of trial syndromes and returns as soon as any one is successfully decoded as GPU can naturally decode multiple syndromes in parallel. We also observe that the BP-SF decoder exhibits a slightly higher minimum runtime (approximately 0.1 ms), which may be attributed to the use of a wrapper function around the CUDA-Q decoder, which introduces additional parameter I/O and initialization overhead.
Figure 16: Distribution of single-syndrome decoding times for GPU decoders when physical error rate is 0.003.
[…]
VII Conclusion and Future Work
@@
We introduced a fully parallelizable decoder based on belief propagation (BP). By leveraging speculative decoding and bit-flipping strategies guided by BP oscillation statistics, our proposed BP-SF method achieves logical error rates comparable to BP-OSD, while significantly reducing computational complexity and avoiding costly Gaussian elimination. Extensive simulations show that our decoder performs exceptionally well under the code-capacity noise model across a range of bivariate bicycle codes. Under the more realistic circuit-level noise model, the decoder still delivers reasonable performance, though it requires a larger number of decoding trials to achieve comparable accuracy. We also want to notice that throughout the paper, we adopt the widely used min-sum variant of BP because of its simplicity and computational efficiency. Nonetheless, our approach could potentially benefit from incorporating more advanced BP-based techniques as long as their convergence is also affected by oscillating bits.
@@
In future work, we aim to better understand the challenges posed by circuit-level noise and explore targeted improvements, such as more effective candidate selection, improved trial vector sampling strategies, efficient decoder implementation, and enhancements to the inner BP decoder, in order to further improve decoding performance in practical fault-tolerant quantum computing systems.
@@ Acknowledgment
The authors would like to thank John Stack for providing source code and for discussions on circuit-level noise simulation, and Timo Hillmann for valuable discussions regarding the BP-OSD decoder. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830. This work was also supported in part by NSF awards OSI-2410675, PHY-1818914, PHY-2325080, MPS-2120757, CISE-2217020, and CISE-2316201 as well as DOE DE-SC0025384.
-A Code Construction
Figure 17: (a-b): The logical error rates of different codes under code capacity error model. For the $\llbracket 72,12,6\rrbracket$ $B$ code, BP and BP-OSD have the same error when we set the seed to the same. BP-SF have $w_{max}=1$, and $|\Phi|=4,7,6,13$ for $\llbracket 72,12,6\rrbracket,\llbracket 144,12,12\rrbracket,\llbracket 126,12,10\rrbracket,\llbracket 254,28\rrbracket$ codes, respectively. (c): Error rate comparison on the $\llbracket 72,12,6\rrbracket$ under circuit-level noise model.
In this section, we provide the construction parameters for the qLDPC codes referenced throughout the paper. The construction and circuit-level implementation of SHYPS code can be found in [16].
-A1 GB codes
@@
$S_{l}=I_{l}>>1,$ (12)
where “>>>>” denotes the right cyclic shift for each row in the matrix, and $I_{l}$ be the identity matrix of size $l$. For example,
$S_{3}=I_{3}>>1=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}$. (13)
Let $x=S_{l}$ here. The GB code can be defined by two polynomials, $a(x)$ and $b(x)$. These two polynomials can be represented by two matrices $A,B$ naturally, and the parity check matrices for the GB code are defined as
$\begin{split}H_{X}&=[A|B]\\ H_{Z}&=[B^{T}|A^{T}].\end{split}$ (14)
The $\llbracket 254,28\rrbracket$ GB code used in this paper can be constructed by $l=127$, $a(x)=1+x^{15}+x^{20}+x^{28}+x^{66}$, $b(x)=1+x^{58}+x^{59}+x^{100}+x^{121}$ as proposed in [18].
-A2 BB Codes
BB codes are constructed similarly to GB codes but with two variables. Let $x=S_{l}\otimes I_{m}$ and $y=I_{l}\otimes S_{m}$, where $\otimes$ denote Kronecker product, the BB codes can also be defined by two polynomials, $A=a(x,y)$ and $B=b(x,y)$. The parity check matrices for BB codes are defined similarly to those in GB codes. The BB codes we used in the paper were proposed in [2]. And the polynomials are shown in Table II.
TABLE II: BB Codes Used in Simulations
@@
-A3 Coprime-BB Codes
The coprime-BB code we used originates from [33]. Let $\pi=xy$, where $x$ and $y$ are defined the same as in BB codes. The coprime-BB code we tested can be constructed as shown in Table III.
TABLE III: Coprime-BB Codes Used in Simulations
@@
This subsection presents several codes that demonstrate good performance under BP decoding, along with their corresponding logical error rates. Since both BP-OSD and the proposed BP-SF decoder act as post-processing techniques, they are only invoked when the BP decoder fails to converge. As a result, for the codes shown below, where BP alone achieves high success rates, we observe similar overall performance across all decoding strategies.
@@
Fig. 17(a-c) shows the performance of various decoders on several codes under the code-capacity and circuit level noise model. In these cases, the baseline BP decoder already achieves logical error rates comparable to those of the BP-OSD decoder, and post-processing yields only marginal improvements.
Appendix A Artifact Appendix
A-A Abstract
This artifact contains the source code for the BP-SF decoder, a parallelized belief-propagation decoder for quantum LDPC codes, as described in the paper “Fully Parallelized BP Decoding for Quantum LDPC Codes Can Outperform BP-OSD”. It includes a custom Cython-based implementation of the Belief Propagation (BP) decoder, scripts for running circuit-level memory experiments using stim, and benchmarking tools to compare performance against standard BP, BP-OSD, and CUDA-Q qLDPC decoders. The artifact allows researchers to reproduce the Logical Error Rate (LER) and decoding speed results presented in the paper.
A-B Artifact check-list (meta-information)
  • Algorithm: The code repository contains BP-SF, BP-OSD, Standard BP (Min-sum) algorithms.
  • Program: Python 3.11 scripts with Cython extensions.
  • Compilation: C/C++ compiler required for Cython extension.
  • Data set: Synthetic data generated on-the-fly using stim circuit simulations.
  • Run-time environment: Linux; Python 3.11-13 (newer versions may work but are untested)
  • Hardware: Standard CPU for functional tests. NVIDIA GPU with CUDA support recommended for cudaq-qec comparisons and GPU estimation experiments.
  • Metrics: Logical Error Rate per Round (LER/Round), average decoding time per sample (ms).
  • Output: Text logs in data/ directory and Jupyter Notebook for plotting.
  • How much disk space required (approximately)?: $\sim$ 200 KBytes.
  • How much time is needed to prepare workflow (approximately)?: $\sim$10 minutes (install dependencies).
  • How much time is needed to complete experiments (approximately)?: Days on a 16-core machine.
  • Publicly available?: Yes. https://github.com/Dies-Irae/BP-SF.
  • Code licenses (if publicly available)?: MIT license.
  • Data licenses (if publicly available)?: Not applicable.
  • Workflow automation framework used?: Not applicable.
  • Archived (provide DOI)?: Not available
A-C Description
A-C1 How to access
The code repository is available at https://github.com/Dies-Irae/BP-SF.
A-C2 Hardware dependencies
The core BP-SF decoder runs on standard CPUs. To reproduce the GPU comparison baselines (using cudaq-qec), an NVIDIA GPU with CUDA support is required.
A-C3 Software dependencies
  • OS: Linux.
  • Python: Version 3.11.
  • Libraries: numpy, scipy, stim, cython, setuptools, matplotlib.
  • Optional (for baselines): ldpc, cudaq-qec.
A-D Installation
  1. 1.
    Clone the repository:
    git clone git@github.com:Dies-Irae/BP-SF.git
    cd BP-SF
  2. 2.
    Create and activate a virtual environment:
    python -m venv .venv
    source .venv/bin/activate
  3. 3.
    Install Python dependencies:
    pip install cython numpy scipy
    pip install stim setuptools matplotlib
    # Optional: For baselines
    pip install cudaq-qec ldpc
  4. 4.
    Build the local Cython extension:
    export PYTHONPATH=./src_python:\$PYTHONPATH
    cd minimal_bp_decoder
    python setup.py build_ext --inplace
    cd ..
A-E Experiment workflow
The artifact provides shell scripts to automate the experiments. Results are saved to the data/ directory.
A-E1 Functional Tests (LER)
To evaluate the logical error rate of the BP-SF decoder compared to BP-OSD:
mkdir data
# Run BP-SF circuit-level simulation
sh bpsf_circ.sh
# Run BP-OSD circuit-level simulation
# (requires ldpc)
sh bposd_circ.sh
A-E2 Performance Benchmarks (Timing)
To measure and compare decoding speeds between BP-SF, BP, BP-OSD, and CUDAQ:
# Run speed benchmarks
sh time.sh
Note: The time.sh script runs benchmarks for varying physical error rates and decoders. Ensure cudaq-qec is installed for the GPU baseline.
A-F Evaluation and expected results
After running the experiments, the results will be stored as text files in the data/ folder (e.g., time_cudaq.txt, *bb_test_*.txt).
A-F1 Plotting
Open the Jupyter notebook plots.ipynb and execute the cells to generate to visualize the results.
  • LER Plots: Comparison of Logical Error Rate vs. Physical Error Rate for BP-SF and BP-OSD (corresponding to Fig. 7-10 in the paper).
  • Timing Plots: Comparison of average decoding time per sample (corresponding to Fig. 14 in the paper).